Dean's and April's ages have a sum of 55 years.
In two years Dean will ne twice as old as April was 12 years ago. Find their present ages

Answer by lwsshak3(11628)

Dean's and April's ages have a sum of 55 years.
In two years Dean will be twice as old as April was 12 years ago. Find their present ages
***
let x=Dean's age
55-x=April's age
..
12 yrs ago:
55-x-12=43-x=April's age
..
In 2 yrs:
x+2=2(43-x)
x+2=86-2x
3x=84
x=28
55-x=27
Dean's age=28
April's age=27